OpenStudy (anonymous):

If cos (A) = 1/2 with A in QIV, find sec (A/2).

OpenStudy (anonymous):

Cos(A) = 1/2 means A = 60 degrees. But, since it is in fourth quadrant, it means A = 270+60 degrees = 330 degrees. So, A/2 = 330/2 = 165 So, sec(A/2) = sec(165) = -1.035

OpenStudy (anonymous):

or, if you want the exact value use the half-angle formula for cosine and take the reciprocal...

OpenStudy (anonymous):

So, what would be the exact angle? I'm sorry, I'm just confusing myself.

OpenStudy (anonymous):

no, I mean the exact value for sec(165 degrees)...

OpenStudy (anonymous):

that is, if the exact value is required...

OpenStudy (anonymous):

Right... So, how would I write the equation? I only have three kinds I learned about, and they are sin, cos, and tan. I don't know how to find sec from cos, though I know they are opposite reciprocals. And, yes. We do have to have the exact value.

OpenStudy (anonymous):

sec = 1/cos

OpenStudy (anonymous):

ok. so you're familiar with the half-angle formula?

OpenStudy (anonymous):

secant is reciprocal of cosine.

OpenStudy (anonymous):

In the above, I calculated cos(165) and did its inverse to figure out sec(165).

OpenStudy (anonymous):

Yes, dpalnc. I just don't know how to use it very well.

OpenStudy (anonymous):

here's the half-angle formula for cosine: